# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

### v1

@@ 4.2 Generative AI-Digital Twin Integration
Figure 5: Technical strategies for AI-Digital Twin integration: Generative (Gen)-AI models deployment in different digital twin architecture.
@@
Through the evolution history of the urban digital twins, a few software architectures emerge as particularly adept at fostering sophisticated, scalable, and efficient digital representations of urban landscapes. Among these, monolithic architecture, microservices, and edge-computing architecture are pivotal (Ferko et al., 2022). Monolithic architecture lays the groundwork by facilitating interoperable services that communicate across a network and is crucial for developing basic cyberinfrastructure for amalgamating diverse simulation systems and data sources using the traditional server-client paradigm (Gos and Zabierowski, 2020). Evolving from monolithic architecture, microservices architecture advances this model by organizing applications into collections of loosely coupled services that are often adopted in large-scale cloud-based applications. This structure promotes agile development and deployment of complex cyberinfrastructure systems, enabling seamless updates and scalability for specific elements (e.g., simulation models for specific urban subsystems) of an urban digital twin without impacting the overall framework (Ponce et al., 2019; Xu et al., 2023b). Meanwhile, Edge computing introduces a move towards decentralized and distributed data processing, allocating computational tasks nearer to the data origins at the network’s edge nodes (Yu et al., 2017). Examples of these edge nodes are often edge servers that are directly connected to sensors and mobile IoT devices for sensing various types of urban data (Xu et al., 2022; Marjanović et al., 2018). This shift is vital for urban digital twin development, as it allows for real-time data analysis and immediate response critical for the effective management of dynamic urban activities. Collectively, these architectural strategies create a formidable foundation for constructing urban digital twins, delivering the adaptability, scalability, and immediacy required to simulate and manage complex urban systems accurately.
@@
To effectively integrate generative AI-powered applications within urban digital twins—often manifested through web-based platforms, cyberinfrastructures, and cloud computing services for enhanced smart city functionalities—we propose several technical strategies (as demonstrated in Figure 5). These strategies are designed to bolster scalability, performance, and precision in the AI-digital twin synergy. Tailored specifically to accommodate urban digital twins, these approaches are aligned with distinct architectural patterns. This ensures that each strategy not only addresses the unique challenges of integrating AI into these complex systems but also leverages the inherent strengths of the chosen architectural framework. By doing so, our approach facilitates the creation of dynamic, responsive urban digital twins that can evolve with the changing needs of smart city ecosystems.
Generative AI as a Service:
@@
To integrate AI models into digital twins that adopt monolithic architecture and microservices architecture, a reliable practice would be hosting these models as web services following the paradigm of the service-oriented architecture (Schnicke et al., 2020). Currently, many cloud computing platforms provide the foundational infrastructure needed for deploying ML and DL models as web services at scale, offering flexible resources and cyber-deliver capability that can dynamically adjust to the computational demands of generative AI. Utilizing the recent containerization technologies, such as Docker, and orchestration tools like Kubernetes can further enhance scalability and manageability by allowing AI components to be deployed as microservices. This modular approach facilitates the seamless integration of generative AI into existing urban digital twin frameworks, ensuring that complex data-driven simulations and data analyses run efficiently across distributed computing environments.
@@
For urban digital twin applications that involve real-time data streaming and simulation capabilities, deploying generative AI models on the edge nodes can reduce latency and bandwidth use by processing data closer to the data source (e.g., sensors and IoT devices), which is crucial for real-time applications like traffic flow optimization, building and grid operations, or emergency response simulations within urban digital twins. Generative AI models can be readily deployed on edge servers or data loggers that are directly connected to urban sensors for processing and distributing raw data. This practice is complemented by the use of data streaming and real-time analytics platforms, such as Apache Kafka and Apache Spark, which can handle the high-velocity data feeds sensed from urban environments and provide the necessary processing power to augment and enrich urban data for supporting enhanced situational awareness and simulation-based predictive analytics in real-time.
GPT-powered Autonomous System:
The integration of generative pre-trained transformer (GPT) models into software tools is becoming increasingly popular, supporting research across various scientific domains. Recent studies have demonstrated the potential of leveraging these models to develop autonomous systems and AI agents (Li and Ning, 2023; Vemprala et al., 2023). Such advancements aid in research and decision-making processes, addressing complex scientific challenges by combining GPT APIs with domain-specific data and simulation models. By employing Python or JavaScript APIs provided by GPT, researchers can seamlessly integrate generative AI capabilities into urban digital twins’ various components. This integration facilitates data integration, augmentation, and enrichment, enhancing AI-assisted urban design, education, decision-making, and the interpretation of intricate urban data and simulation models. This approach not only streamlines the incorporation of advanced AI technologies into urban planning and management but also opens new avenues for innovative, data-driven solutions to urban challenges.
4.3 Challenges and Limitations
Although numerous generative AI approaches have been developed to address challenges across different sectors of smart cities, significant obstacles persist. These challenges are often linked to the limitations of the generative AI models themselves or stem from technical issues related to integrating AI with urban digital twins.
@@ 4.3 Challenges and Limitations
The quality of AI-generated data is influenced by several complex factors, including the absence of objective evaluation criteria, challenges in detecting mode collapse, the lack of benchmark data, and issues arising from imbalanced data (Saxena and Cao, 2021). Furthermore, there is a notable lack of consensus regarding assessment indicators for various types of AI-generated data, such as time series and spatial data. This ambiguity in evaluation standards can be attributed partly to the limited number of published studies in this area, which restricts the development of universally accepted benchmarks and metrics.
Computing, Data Storage, and Latency:
In the context of smart cities and urban digital twin applications, generative AI introduces substantial computational challenges, necessitating high processing power to facilitate complex simulations of urban systems and real-time data analytics. Moreover, these applications demand rapid and efficient data storage solutions to handle the extensive data volumes generated and processed, encompassing everything from sensor data to detailed environmental models. The seamless integration of these technologies requires robust, scalable computing infrastructures and sophisticated data storage systems, designed to support the relentless and intensive data streams inherent to smart urban environments.
Network Latency:
The network latency associated with the data exchange for sending the data to generative AI models must be minimized to support real-time smart city functionalities that perform time-sensitive analytics and cyber-physical controls, such as traffic management optimization and incident detection. Addressing these demands often entails equipping urban digital twins with ample communication bandwidth and computational resources to process these AI models in real time, ensuring efficient and effective smart city operations. In fact, generative AI models, such as the VAE, can also be used for data compression, potentially addressing the network latency issues.
@@ 5 Conclusion
This survey investigates and summarizes the recent progress on popular generative AI models and their applications in the smart city domain. To enhance the reliability, performance, and efficiency of urban planning and management, the augmentation and synthesis of urban data based on previously collected and simulated data receive a lot of attention in the transportation, urban mobility, energy systems, infrastructure, and urban design sectors. Generative AI models support these sectors’ urban research and management efforts through their excel capabilities for data augmentation, synthetic data and scenario generation, 3D city modeling, and generative urban design. With the advent of urban digital twin technologies, we envision an integration between the generative AI models and cognitive digital twins for a more reliable, autonomous, scalable, and cost-effective management and digital transformation of complex urban systems. Based on a structured review of the previous generative AI application in the smart city sector, this survey outlines 6 promising opportunities for the proposed generative AI-enabled urban digital twins and outlines 3 technical strategies to facilitate the generative AI-digital twin integration within different software engineering architectures. Finally, the survey closes with a discussion of 4 potential challenges that may present in the AI-digital twin integration processes. We believe the integration of generative AI models could significantly revolutionize the existing urban digital twin applications towards more intelligent, reliable, and autonomous systems for supporting smart city development and operations.

### v2

@@ 4.1.4 Human-AI Partnership
Figure 5: Technical strategies for AI-Digital Twin integration: Generative (GenAI) models deployment in different digital twin architecture.
By integrating GenAI models into urban digital twins, we can create opportunities for generating more inclusive, adaptable, and responsive solutions to the evolving needs and values of various urban management participants. These generative approaches not only enhances the quality and sustainability of urban environments but also empowers communities to play an active role in shaping their future by building transparent and trustworthy human-AI partnership. This approach could provide efficient solution to address challenge C6 discussed in Section 1.2.
4.2 GenAI-Digital Twin Integration: Technical Strategies
@@
This subsection explores potential technical implementation strategies for integrating popular GenAI models into existing urban digital twin applications. Through the evolution history of the urban digital twins, a few software architectures emerge as particularly adept at fostering sophisticated, scalable, and efficient digital representations of urban landscapes. Among these, monolithic architecture, microservices, and edge-computing architecture are pivotal (Ferko et al., 2022). Monolithic architecture lays the groundwork by facilitating interoperable services that communicate across a network and is crucial for developing basic cyberinfrastructure for amalgamating diverse simulation systems and data sources using the traditional server-client paradigm (Gos and Zabierowski, 2020). Evolving from monolithic architecture, microservices architecture advances this model by organizing applications into collections of loosely coupled services that are often adopted in large-scale cloud-based applications. This structure promotes agile development and deployment of complex cyberinfrastructure systems, enabling seamless updates and scalability for specific elements (e.g., simulation models for specific urban subsystems) of an urban digital twin without impacting the overall framework (Ponce et al., 2019; Xu et al., 2023b). Meanwhile, Edge computing introduces a move towards decentralized and distributed data processing, allocating computational tasks nearer to the data origins at the network’s edge nodes (Yu et al., 2017). Examples of these edge nodes are often edge servers that are directly connected to sensors and mobile IoT devices for sensing various types of urban data (Xu et al., 2022b; Marjanović et al., 2018). This shift is vital for urban digital twin development, as it allows for real-time data analysis and immediate response critical for the effective management of dynamic urban activities. Collectively, these architectural strategies create a formidable foundation for constructing urban digital twins, delivering the adaptability, scalability, and immediacy required to simulate and manage complex urban systems accurately.
@@
To effectively integrate genAI-powered applications within urban digital twins—often manifested through web-based platforms, cyberinfrastructures, and cloud computing services for enhanced smart city functionalities—we propose several technical strategies (as demonstrated in Figure 5). These strategies are designed to bolster scalability, performance, and precision in the AI-digital twin synergy. Tailored specifically to accommodate urban digital twins, these approaches are aligned with distinct architectural patterns. This ensures that each strategy not only addresses the unique challenges of integrating GenAI models into these digital twins’ complex systems but also leverages the inherent strengths of the chosen architectural framework. By doing so, our approach facilitates the creation of dynamic modular, and maintainable urban digital twins that can evolve with the changing needs of smart city ecosystems. From the technical perspective, we propose the following strategies:
GenAI as a Service:
@@
To integrate AI models into digital twins that adopt monolithic architecture and microservices architecture, a reliable practice would be hosting these models as web services following the paradigm of the service-oriented architecture (Schnicke et al., 2020). Currently, many cloud computing platforms provide the foundational infrastructure needed for deploying ML and DL models as web services at scale, offering flexible resources and cyber-deliver capability that can dynamically adjust to the computational demands of GenAI. Utilizing the recent containerization technologies, such as Docker, and orchestration tools like Kubernetes can further enhance scalability and manageability by allowing AI components to be deployed as microservices. This modular approach facilitates the seamless integration of GenAI models into existing urban digital twin frameworks, ensuring that complex simulations and data analyses run efficiently across distributed computing environments.
@@
For urban digital twin applications that involve real-time data streaming and simulation capabilities, deploying GenAI models on the edge nodes can reduce latency and bandwidth use by processing data closer to the data source (e.g., sensors and IoT devices), which is crucial for real-time applications like traffic flow optimization, building and grid operations, or emergency response simulations within urban digital twins. GenAI models can be readily deployed on edge servers or data loggers that are directly connected to urban sensors for processing and distributing raw data. This practice is complemented by the use of data streaming and real-time analytics platforms, such as Apache Kafka and Apache Spark, which can handle the high-velocity data feeds sensed from urban environments and provide the necessary processing power to augment and enrich urban data for supporting enhanced situational awareness and simulation-based predictive analytics in real-time.
GPT Prompt Base and Templates for Smart Cities:
The emerging concept of GPT Prompt Template Base involves developing structured templates and foundational setups that guide Generative Pre-trained Transformers (GPT) in generating accurate and contextually relevant outputs. These templates include predefined patterns and placeholders that can be filled with specific information, ensuring consistency and efficiency in prompt creation. By facilitating the autonomous generation of data and code, GPT Prompt Template Bases can significantly streamline the development of urban digital twins. They enable domain experts, such as urban planners and researchers, to easily generate complex data analyses, predictive models, and software components for an urban digital twin without extensive programming expertise. This is particularly important for urban sciences and management, where the integration of diverse datasets and the need for rapid, adaptive decision-making are critical. Creating robust GPT Prompt Template Bases for these fields ensures that AI-generated content is reliable, relevant, and aligned with the specific needs of urban development and management, ultimately enhancing the effectiveness and scalability of urban digital twin applications.
We envision that through the proposed AI-digital twin integration, next-generation urban digital twins could provide more intelligent, intuitive, and straightforward interfaces to deliver personalized smart city services to various IoT-connected devices. These devices include mobile devices (e.g., smartphones, tablets, laptops, and desktops), which host CDT software (e.g., web apps or native apps) as AI agents to directly deliver smart city services to users. Additionally, smart infrastructure can sense real-time urban data and receive cyber-physical controls from the urban digital twins to optimize real-world urban systems. We present several potential use cases that would benefit from the proposed GenAI-digital twin integration: (a) A smart city service delivered through an AI agent that allows city residents to request the generation and visualization of new park designs or building renovation plans, visualized through a mobile augmented reality environment rendered via the user’s smartphone camera view. (b) A smart city service deployed on edge devices to improve the density and quality of traffic data collected through an IoT-connected Radar Detection System (RDS) sensor. This service delivers the enhanced and enriched RDS data to the urban digital twin’s computing environment, providing urban planners with more accurate situational awareness of traffic flow and lane occupancy on urban roads.
5 Challenges and Limitations
Although numerous GenAI applications have been developed to address challenges across various sectors of smart cities, significant obstacles persist. These challenges often arise from the limitations of GenAI models from both data science and computational perspectives (Duan et al., 2019), as well as from socio-ethical issues related to the AI-digital twin integration (Filipenko et al., 2024). In this regard, we categorize these challenges based on their nature in the following subsections and provide a detailed discussion of each challenge, along with potential solutions.
5.1 Data Science Challenges
The data science challenges of GenAI applications refer to the significant demands on computational resources required for training, deploying, and maintaining these models, including high processing power, extensive memory, and energy consumption, which can impact both cost and environmental sustainability (Wang et al., 2023d).
@@ 5.1 Data Science Challenges
Computing, Data Storage, and Latency:
In the context of smart cities and urban digital twin applications, GenAI introduces substantial computational challenges, necessitating high processing power to facilitate complex simulations of urban systems and real-time data analytics (Karapantelakis et al., 2024). Moreover, these applications demand rapid and efficient data storage solutions to handle the extensive data volumes generated and processed, encompassing everything from sensor data to detailed environmental models. The seamless integration of these GenAI models requires robust, scalable computing infrastructures and sophisticated data storage systems, designed to support the relentless and intensive data streams inherent to smart urban environments (Wang et al., 2023d).
@@
The quality of AIGC is influenced by several complex factors, including the absence of objective evaluation criteria, challenges in detecting mode collapse, the lack of benchmark data, and issues arising from imbalanced data (Saxena and Cao, 2021). Furthermore, there is a notable lack of consensus regarding assessment indicators for various types of AIGC, such as time series and spatial data (Lin et al., 2023b; Debnath et al., 2021). This ambiguity in evaluation standards can be attributed partly to the limited number of published studies in this area, which restricts the development of universally accepted benchmarks and metrics (Bandi et al., 2023).
Based on the Moore’s Law and Beyond, continued improvements in semiconductor technology and the advent of novel computing architectures (quantum computing, neuromorphic computing) are expected to significantly enhance computational power (Hiremane, 2005). The recent development of AI-specific hardware, such as GPUs, TPUs, and custom AI accelerators, optimizes the performance of AI models and simulations, making the large-scale integration of GenAI and digital twins more feasible and affordable (Mody, 2016).
5.2 Ethical and Societal Challenges
Building Urban digital twins and smart cities often involve a range of sociotechnical aspects that encompass both the social and technical dimensions of these technologies. From the social perspective, these dimensions include governance and policy, ethical considerations, social equity, and cultural and behavioral changes. Within these applications, GenAI models pose significant ethical and societal challenges, especially in terms of bias and fairness, as highlighted by Ferrara (2024). Additionally, these models raise concerns about transparency and accountability.(Kenthapadi et al., 2023).
5.2.1 Bias and Fairness
GenAI models, such as GAN and GPT, are trained on large datasets that may reflect the biases present in the real world (Grossman et al., 2023). These biases can manifest in various ways, leading to unfair or discriminatory outcomes. Addressing bias in AI is essential for maintaining public trust and acceptance (Schwartz et al., 2022). Bias in AI models primarily originates from two sources: training data and algorithmic design. Training data can carry societal biases related to race, gender, age, and other factors, which the AI model can then replicate and even amplify (Mehrabi et al., 2021). In the context of urban digital twins, societal biases.
Even with unbiased data, algorithms themselves can introduce bias through their design and implementation. The impacts of these biases are profound, leading to discriminatory outcomes where certain groups are favored over others, resulting in unfair treatment (Ferrara, 2023). This not only affects individuals adversely but also erodes public trust in AI technologies, potentially hindering their broader adoption. To mitigate these biases, several strategies are essential. Using diverse datasets that accurately represent various demographics and target user groups can help reduce inherent biases (Nazer et al., 2023).
Employing bias detection tools during the development and deployment phases allows for the identification and mitigation of biases early on. Additionally, conducting regular audits of AI systems ensures ongoing fairness and the ability to address emerging biases effectively. These steps are critical for creating AI models that are equitable, trustworthy, and widely accepted.
5.2.2 Transparency and Accountability in AI Systems
The complexity and opacity of AI models, often referred to as the ”black-box” nature, pose significant challenges for understanding and explaining their decisions (Von Eschenbach, 2021). Ensuring transparency and accountability is crucial for building trust and meeting regulatory requirements. Black-box models suffer from a lack of interpretability, making it difficult for humans to understand and explain their decision-making processes (Hassija et al., 2024). This lack of transparency can lead to issues with regulatory compliance, as many industries require clear and explainable AI systems. Transparency is essential for building stakeholder trust, as users and stakeholders are more likely to trust AI systems that are clear and understandable (Balasubramaniam et al., 2022). Additionally, transparency is crucial for the ethical use of AI, ensuring that these systems are used responsibly and without causing harm.
To ensure accountability, the development of explainable AI (XAI) is vital, as it provides explanations for AI decisions, enhancing transparency (Arrieta et al., 2020). Thorough documentation of AI systems, including their development, training data, and decision-making processes, is also necessary. Furthermore, establishing governance frameworks to oversee the ethical use and accountability of AI systems is essential for maintaining transparency and trust in AI technologies (Díaz-Rodríguez et al., 2023).
@@ 6 Conclusion
This survey investigates and summarizes the recent progress on popular GenAI models and their applications in the smart city domain. To enhance the reliability, performance, and efficiency of urban planning and management, the augmentation and synthesis of urban data based on previously collected and simulated data receive a lot of attention in the transportation, urban mobility, energy systems, infrastructure, and urban planning sectors. GenAI models support these sectors’ urban research and management efforts through their excel capabilities for data augmentation, synthetic data and scenario generation, 3D city modeling, and generative urban design. With the advent of urban digital twin technologies, we envision an integration between the GenAI models and cognitive digital twins for a more reliable, autonomous, scalable, cost-effective, and participatory management and digital transformation of complex urban systems. Based on a structured review of the previous GenAI application in the smart city sector, this survey outlines 3 promising directions for the proposed GenAI-enabled urban digital twins and outlines 3 technical strategies to facilitate the GenAI-digital twin integration within different software engineering architectures. Finally, the survey closes with a discussion of potential challenges that may present in the GenAI-digital twin integration processes in the context of computing and data science limitations, and ethical and societal concerns. We believe the integration of GenAI models could significantly revolutionize the existing urban digital twin applications towards more intelligent, reliable, and autonomous systems for supporting smart city development and operations.